OpenStudy (moose213):

Help with these 3 questions please? =) 1. Martin needs 3.25 gallons of red paint for his small barn and 1.25 gallons of paint for his shed. He spends $81 on paint. How much does the paint cost per gallon? $18.00 $24.54 $27.00 $40.50 2. Margo spends 0.25 of her paycheck on a new dress, 1/8 of her paycheck on shoes, and $24 on a birthday present. If she spent $60 altogether, how much was Margo’s paycheck? $96 $160 $216 $504 3. What is the first step in solving the equation 3.5n + 6.4 = 42.5

OpenStudy (moose213):

Add 6.4 to each side of the equation. Subtract 6.4 from each side of the equation. Multiply each side of the equation by 3.5. Divide each side of the equation by 6.4.

OpenStudy (jjcoolja12):

#3 Let's solve your equation step-by-step. 3.5n+6.4=42.5 Step 1: Subtract 6.4 from both sides. 3.5n+6.4−6.4=42.5−6.4 3.5n=36.1 Step 2: Divide both sides by 3.5. 3.5n 3.5 = 36.1 3.5 n=10.314286 Answer: n=10.314286

OpenStudy (moose213):

What is the first step in solving the equation 3.5n + 6.4 = 42.5 Add 6.4 to each side of the equation. Subtract 6.4 from each side of the equation. Multiply each side of the equation by 3.5. Divide each side of the equation by 6.4.

OpenStudy (moose213):

That's #3

OpenStudy (jjcoolja12):

subtract 6.4

OpenStudy (moose213):

#2?

OpenStudy (jjcoolja12):

id say 160

OpenStudy (moose213):

#1?

OpenStudy (jjcoolja12):

c

OpenStudy (moose213):

27 dollars?

OpenStudy (jjcoolja12):

ok i may have did that one wrong

OpenStudy (moose213):

...

OpenStudy (jjcoolja12):

i think its actually a

OpenStudy (moose213):

Ok

OpenStudy (moose213):

What is the value of x? x+ 4/9 = 3/4

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. x+ 4 9 = 3 4 Step 1: Subtract 4/9 from both sides. x+ 4 9 − 4 9 = 3 4 − 4 9 x= 11 36 Answer: x= 11 36

OpenStudy (moose213):

What is the first step in solving the equation 2/3 x + 1/3 x +2 =5?

OpenStudy (moose213):

Subtract 1/3 from each side of the equation. Add 2 to each side of the equation. Combine like terms. Multiply each side of the equation by 5.

OpenStudy (jjcoolja12):

combine like terms

OpenStudy (moose213):

Cooper is measuring ingredients for his special dessert. He needs a total of 70.4 mL of milk and melted butter. The recipe calls for 30.6 mL of butter. How much milk, m, does Cooper need for his recipe?

OpenStudy (moose213):

2.3 mL 39.8 mL 40.2 mL 101 mL

OpenStudy (jjcoolja12):

2.3ml you almost had me there

OpenStudy (moose213):

A recipe calls for a total of 3 2/3cups flour and sugar. If the recipe calls for 1/4 cup of sugar, how much flour is needed? By the way, the 3 in 3 2/3 is a whole number

OpenStudy (jjcoolja12):

0.916667 you may have me on this one

OpenStudy (moose213):

It's alright

OpenStudy (moose213):

Which is the solution to the equation 8.25 +1/4w = 10.75 Round to the nearest hundredth if necessary.

OpenStudy (moose213):

0.63 4.75 10 76

OpenStudy (jjcoolja12):

it is 10 ms pc was going slow

OpenStudy (jjcoolja12):

*my

OpenStudy (moose213):

Last winter Armand had 5/6 of a row of stacked logs. At the end of the winter he had 8/15 of the same row left. How much wood did he burn over the winter?

OpenStudy (moose213):

Last one, I'll send it to you after you answer this one =)

OpenStudy (jjcoolja12):

0.017778 i think

OpenStudy (moose213):

Ok I got 90% thanks I sent them check your mail

OpenStudy (moose213):

Let me know on here if you got them.

OpenStudy (jjcoolja12):

yep i got them sexy

OpenStudy (moose213):

I'll send some of my private parts if you help with 10 more.

OpenStudy (jjcoolja12):

ok!

OpenStudy (moose213):

Give me 5 minutes

OpenStudy (jjcoolja12):

close the question

OpenStudy (moose213):

Ok

OpenStudy (moose213):

Why can't I post them here?

OpenStudy (jjcoolja12):

mods rules

OpenStudy (moose213):

When I post the question just go on there.

OpenStudy (jjcoolja12):

k

OpenStudy (jjcoolja12):

ms.hottie

OpenStudy (moose213):

*wink*